# Token Embeddings Violate the Manifold Hypothesis

### v2

@@ Abstract.
A full understanding of the behavior of a large language model (LLM) requires our understanding of its input token space. If this space differs from our assumptions, our understanding of and conclusions about the LLM will likely be flawed. We elucidate the structure of the token embeddings both empirically and theoretically. We present a novel statistical test assuming that the neighborhood around each token has a relatively flat and smooth structure as the null hypothesis. Failing to reject the null is uninformative, but rejecting it at a specific token $\psi$ implies an irregularity in the token subspace in a $\psi$-neighborhood, $B(\psi)$. The structure assumed in the null is a generalization of a manifold with boundary called a smooth fiber bundle (which can be split into two spatial regimes – small and large radius), so we denote our new hypothesis test as the “fiber bundle hypothesis.” Failure to reject the null hypothesis is uninformative, but rejecting it at $\psi$ indicates a statistically significant irregularity at $B(\psi)$. By running our test over several open-source LLMs, each with unique token embeddings, we find that the null is frequently rejected, and so the evidence suggests that the token subspace is not a fiber bundle and hence also not a manifold. As a consequence of our findings, when an LLM is presented with two semantically equivalent prompts, if one prompt contains a token implicated by our test, the response to that prompt will likely exhibit less stability than the other.
1. Introduction
@@
A core area for Artificial Intelligence (AI) research is to understand the central components of modern learning systems, e.g. large language models (LLMs), assuming certain regularity conditions. Extensive research efforts have been mobilized to tackle the explainability of AI question ([1]), leveraging ideas from computer science ([2, 3, 4]), information theory ([5, 6]), and theoretical physics ([7, 8, 9]), to mathematical insights from high-dimensional probability theory ([10, 11]), and algebraic and differential topology ([12, 13, 14]) either to prove mathematical guarantees or to elucidate model behavior.
@@
In many—if not most—AI research papers, a manifold hypothesis is tacitly assumed, that the data are concentrated near a low curvature manifold without boundary. This assumption drives the research just as an assumption of Gaussianity would drive data modeling. It is important to test this kind of assumption before data modeling. One common strategy to perform such a test is simply “plotting the data” to ascertain the empirical distribution. Plotting high-dimensional data, however, is non-trivial, but there are statistical tests to help us understand the underlying distribution (Mann-Whitney, Kolomogorov-Smirnoff, etc.). In this paper, we provide an analogous test for the regularity of token embeddings.
@@
The first layer of most LLMs is usually comprised of the token embeddings, which are vector representations of tokens (fragments of human language, e.g. parts of words, or numbers). Thus, a good first step to explain the behavior of an LLM is to understand its token subspace where these embedding vectors live. For the remainder of this paper, the term “token”’ used in the context of the token subspace will refer to this embedding. Numerous papers (for instance, see [15, 16, 17, 18, 19, 20, 21]) have pointed to unexpected behaviors exhibited by LLMs that hinge on subtle changes in wording and text layout between apparently similar prompts, suggesting that certain semantically similar tokens may have dramatically different neighborhoods in the token subspace. These differences in neighborhoods correspond to places where the token subspace does not behave stably as it is irregular at such a token. Linguistically, these irregularities may correspond to polysemy or homonyms—tokens with multiple distinct meanings ([22]).
@@
The presence of irregularities imply small changes in language can drive large variation in the token subspace. Moreover, if the token subspace of an LLM is, in fact, singular (not a manifold), irregularities can persist into the output of the LLM, perhaps unavoidably and potentially even because of its architecture. This is because the input and output of LLMs are a fixed finite set of tokens yet the internal architectures of the model assume a structure resembling a Riemannian manifold: a metrizable space that locally preserves inner products. Not accounting for singularities in the token subspace may confound our understanding of the LLM’s behavior.
1.1. Contributions
We present three hypothesis tests that determine for a given token $\psi$, whether its token subspace neighborhood $B(\psi)$ contains an irregularity, i.e. either has significantly high curvature or is singular. The tests are based on Theorem 1 that allows us to identify changes in subspace dimension that are inconsistent with the token subspace being:
  • 1.
    A manifold (without boundary) with a given embedded reach (low reach corresponds to high curvature and vice versa ([23])),
  • 2.
    A fibered manifold with boundary (which we call a fiber bundle for brevity) with the same embedded reach, estimated using small radii, and
  • 3.
    A fiber bundle with the same embedded reach, estimated using large radii.
Figure 1. This figure shows the application of this paper’s manifold hypothesis test to a neighborhood of the token ember in Mistral7B. (The fiber bundle tests are not shown due to space considerations.) Since the latent space of Mistral7B has dimension $\ell=4096$, we have projected the coordinates of each token shown to $3$ using principal components analysis (PCA). The inset shows all of the tokens in Mistral7B (rendered with t-SNE), with a crosshairs showing the location of the zoomed-in region. Each point shown in this figure is a token used by Mistral7B, colored based upon the $p$-value for the manifold hypothesis test. Darker colors correspond to lower $p$-values, which are less consistent with the hypothesis that the token’s neighborhood is a manifold with low curvature. Putative geometric structures are labeled for certain tokens.
@@ 1.1. Contributions
When these tests are rejected, this implies that $\psi$ is less stable under mathematical transformations than other tokens. In Section 3.2, we show the endemic nature of these instabilities by proving that context cannot resolve singularities persistently. And finally, we provide conditions (Theorem 2) for when singularities can propagate into the output of an LLM for generic transformers.
In our work, we provide empirical evidence that the global token embedding obtained from four different LLMs cannot be a manifold due to non-constant local dimension. Because our hypothesis tests require the distances between all pairs of tokens, we need the token embedding vectors. For this reason, we conduct our analysis to four open source models of moderate size – GPT2 ([24]), Llemma7B ([25]), Mistral7B ([26]), and Pythia6.9B ([27]). We emphasize that while this paper demonstrates the manifold and fiber bundle tests on four open source LLMs, our main contribution is primarily a methodology for hypothesis testing on embedded points. In particular, our method can be applied to any point cloud data set; the token embeddings we study here are merely one example.
Figure 2. (Top) Flowchart of our statistical methodology. (Bottom) The distribution of local dimensions estimated near tokens in GPT2, produced by our method as an intermediate step. This histogram shows a mixture with at least three distinct peaks.
1.2. Background and related work
@@
Manifold learning takes its roots in nonlinear dimensionality reduction of high dimensional data into low dimensional representations of smooth manifolds. While these algorithms have played an important role in modern data science especially in domains such as cosmology ([28]), astrophysics ([29, 30]), dynamical systems ([31, 32]), and quantum chemistry ([33]), the main utility of these methods have been for data visualization. For instance, even though the Uniform Manifold Approximation and Projection (UMAP) provided an optimal algorithm for fitting sampled data to a manifold, the authors themselves concede that “if the global structure is of primary interest then UMAP may not be the best choice for dimensionality reduction” ([34]). These manifold learning algorithms will find some smooth structure to fit the data rather than “error out” when the data is not sampled from a manifold as it is not a hypothesis test. While UMAP will fit a manifold to data optimally, for instance, the manifold it finds may have extremely high curvature or dimension. Both of these are undesirable properties in a model, and both implicitly suggest that a manifold may be a poor choice for a representation. To address this concern, [35] presented a test for whether data lie on a manifold with a specified embedded reach. Unfortunately, [35] does not include the values of key constants, which makes the practical usage of this test impossible. Our paper remedies both issues: we construct a practical hypothesis test for manifolds and fiber bundles with a specified embedded reach and implement it as software.
@@
In our test case, the basic assumption is that transformers act on the entire extrinsic latent space containing the token embedding. This is clearly a manifold as it is Euclidean. But the truth is that a transformer in the context of an LLM really only acts on the embedded token subspace. The token input embedding matrix defines this subspace, by specifying where the tokens are located. Such a space is quite unconstrained. There is no a priori reason to suspect it is a manifold. It has already been shown that local neighborhoods of each token have salient topological structure ([13]), and one of the most basic parameters is the dimension near any given token in this space. Higher dimension at a token $\psi$ means that $\psi$ has greater density while lower dimensional tokens have less density ([22]). Intuitively, a densely populated $B(\psi)$ will generically occupy a higher dimension than otherwise.
@@
Assuming that word embeddings yield manifolds, some researchers have used global dimension estimators on token input embeddings and word embeddings ([36, 37]). By using a local (not global) dimension estimator, [38] presented the first (to our knowledge) direct test of whether the token subspace is a manifold for the token input embeddings for several LLMs, though no rigorous hypothesis test was presented. Nevertheless, a strongly negative result was obtained: the subspace of tokens is apparently never a manifold, so global dimension estimators are not reliable.
2. Methods
Our method, summarized in Figures 2 and 3, proposes three related hypotheses: (1) the manifold hypothesis, (2) the fiber bundle hypothesis for small radius, and (3) the fiber bundle hypothesis for larger radius.
Figure 3. (a) Schematic of our three hypothesis tests (b) Application of the tests to the tokens #, ¢, and \$ with GPT2’s token input embedding.
@@ 2. Methods
The key insight is that Theorem 1 asserts that for a fiber bundle, the number of tokens within a ball of a given radius $r$ of a fixed token is a piecewise log-log-linear function of radius, in which the slopes correspond to dimension and decrease as $r$ increases. If there is a statistically significant change in the slope estimates (here, the significance level $\alpha=10^{-3}$), then we reject the manifold hypothesis. If there is a statistically significant increase in slope, then we additionally reject the fiber bundle hypothesis in the corresponding region (small or large radius).
Figure 3(b) demonstrates our method on three tokens used by GPT2. The tokens \$ and # play an important role in many programming languages, whereas ¢ does not. (Note that these tokens were chosen for illustrative purposes. The $p$-value for rejecting the hypotheses is larger than $\alpha=10^{-3}$ used in Section 3.3, so these particular tokens do not appear in any of the tables in the following sections.) The token # does not show any rejections of our hypotheses. While this does not allow one to conclude that the vicinity of # is a manifold or fiber bundle, we can use Theorem 1 to estimate the dimension from the slope of the curve (approximately $53$, once logarithms of the values on both axes are taken). The token ¢ exhibits a rejection of the manifold hypothesis, but not the fiber bundle hypothesis. We can conclude that there is a irregularity at a distance of roughly $3.9$ units of ¢. The curve for \$ exhibits two slope changes at radii $3.8$ and $4.0$. The smaller radius slope change represents a violation of the manifold hypothesis, while the large radius slope change represents a violation of the fiber bundle hypothesis.
3. Results
3.1. Volume versus radius for fiber bundles
@@
The learned token embedding vectors in an LLM each have the same number of elements, which we refer to as $\ell$. Thus, the explicit representation of the token subspace $T$ used by an LLM arises by embedding $T$ within the latent space $\mathbb{R}^{\ell}$ by way of a smooth function $e:T\to\mathbb{R}^{\ell}$. Since embeddings are bijective, each $\psi\in T$ corresponds to a unique element $e(\psi)\in\mathbb{R}^{\ell}$. With a slight abuse of notation, we can write $\psi\in e(T)$.
The fiber bundle model imposes a specific assumption about the structure on $T$, by relating it to another, lower dimensional, manifold $S$, called the base space, via a smooth map $p:T\to S$.
Definition 1.
A fibered manifold is a surjective function $p:T\to S$ such that the Jacobian matrix $J_{\psi}p$ at every point $\psi\in T$ has rank equal to the dimension of $S$. For brevity, we refer to this kind of space simply as a fiber bundle.
If a fiber bundle is the correct representation of the space of tokens $T$, then the probability distribution of tokens given by the volume form $v$ on $T$ will impact the distribution of tokens within $\mathbb{R}^{\ell}$.
@@
Suppose that $T$ is a compact, finite-dimensional Riemannian manifold with boundary [^1] Every point in a manifold with boundary has a neighborhood that is locally homeomorphic to a half-space. As a consequence, manifolds are a special case of manifolds with boundary., with a volume form $v$ satisfying $v(T)<\infty$, and let $p:T\to S$ be a fibered manifold. If $e:T\to\mathbb{R}^{\ell}$ is a smooth embedding with reach $\tau$, then there is a function $\rho:e(T)\to[0,\tau]$ such that if $\psi\in e(T)$, the induced volume $(e_{*}v)$ in $\mathbb{R}^{\ell}$ satisfies
(1) $(e_{*}v)\left(B_{r}(\psi\right)=\begin{cases}O\left(r^{\dim T}\right)&\text{if}0\leq r\leq\rho(\psi),\\ (e_{*}v)\left(B_{\rho(\psi)}(\psi)\right)+O\left((r-\rho(\psi))^{\dim S}\right)&\text{if }\rho(\psi)\leq r,\\ \end{cases}$
@@
See Section A.4
Theorem 1 characterizes the resulting probability distribution using parameters (the exponents in Equation (1)) that can be estimated from the token input embedding. These parameters are bounded by the dimensions of $S$ and $T$.
In short, if we take the logarithm of both the volume and radius in Equation (1), we obtain a piecewise linear curve, in which the slopes are given by $\dim S<\dim T$, and the slopes must decrease through a discontinuity as the radius increases, as shown in Figure 3(a).
3.2. Context does not persistently resolve singularities
Given a context window of $w-1$ tokens, an LLM uses its learned parameters to predict the next ($w$-th) token. This gets added to the context window and the process continues. (See Section A.1 for details.) In this section, we provide a mathematical argument showing that the use of context is insufficient to resolve singularities present in the token subspace for a generic transformer in a persistent manner. Again, denote the token subspace by $T$, noting that Section 3.3 shows that $T$ is unlikely to be a manifold for typical LLMs. Let us begin with a lemma:
Lemma 1.

### v3

@@ Abstract
A full understanding of the behavior of a large language model (LLM) requires our grasp of its input token space. If this space differs from our assumptions, our comprehension of and conclusions about the LLM will likely be flawed. We elucidate the structure of the token embeddings both empirically and theoretically. We present a novel statistical test assuming that the neighborhood around each token has a relatively flat and smooth structure as the null hypothesis. Failing to reject the null is uninformative, but rejecting it at a specific token $\psi$ implies an irregularity in the token subspace in a $\psi$-neighborhood, $B(\psi)$. The structure assumed in the null is a generalization of a manifold with boundary called a smooth fiber bundle (which can be split into two spatial regimes – small and large radius), so we denote our new hypothesis test as the “fiber bundle hypothesis.” By running our test over several open-source LLMs, each with unique token embeddings, we find that the null is frequently rejected, and so the evidence suggests that the token subspace is not a fiber bundle and hence also not a manifold. As a consequence of our findings, when an LLM is presented with two semantically equivalent prompts, if one prompt contains a token implicated by our test, the response to that prompt will likely exhibit less stability than the other.
1 Introduction
@@
A core area for Artificial Intelligence (AI) research is to understand the central components of modern learning systems, e.g. large language models (LLMs), assuming certain regularity conditions. Extensive research efforts have been mobilized to tackle the explainability of AI question (Gunning et al. (2019)), leveraging ideas from computer science (Gordon (2024); Wu et al. (2022); Traylor et al. (2021)), information theory (Tan et al. (2024); Wang et al. (2024)), and theoretical physics (Chen et al. (2025); Geshkovski et al. (2023); Lu et al. (2019)), to mathematical insights from high-dimensional probability theory (Huang et al. (2024); Asher et al. (2023)), and algebraic and differential topology (Bradley and Vigneaux (2025); Rathore et al. (2023); Bradley et al. (2022)) either to prove mathematical guarantees or to elucidate model behavior.
@@
In many—if not most—AI research papers, a manifold hypothesis is tacitly assumed, that the data are concentrated near a low curvature manifold without boundary. This assumption drives the research just as an assumption of Gaussianity drives classical approaches to data modeling. For instance, the Uniform Manifold Approximation and Projection (UMAP) algorithm (McInnes et al. (2018)) explicitly assumes that data are sampled near a low dimensional manifold and cautions the reader that the algorithm may not be reliable if this assumption is violated. Yet many researchers make inferences from UMAP, rarely checking that this assumption holds, and even arguing it is unnecessary. This argument has been routinely rejected in the literature (Jeon et al. (2025); Chari and Pachter (2023)), and so it is important to test this kind of assumption before data modeling. One common testing strategy is simply “plotting the data” to ascertain the empirical distribution. Plotting high-dimensional data, however, is non-trivial, but there are statistical tests to help us understand the underlying distribution (Mann-Whitney, Kolomogorov-Smirnoff, etc.). In this paper, we provide an analogous test for the regularity of token embeddings.
@@
The first layer of most LLMs is usually comprised of the token embeddings, which are vector representations of tokens (fragments of texts). Thus, a good first step to explain the behavior of an LLM is to understand its initial token subspace where these embedding vectors lie. For the remainder of this paper, the term “token”’ used in the context of the token subspace will refer to this embedding. Numerous papers (for instance, see Verma et al. (2024); He et al. (2024); Vargas et al. ((to appear); Li et al. (2024); Battle and Gollapudi (2024); Valmeekam et al. (2023); Sclar et al. (2023); Chao et al. (2023)) have pointed to unexpected behaviors exhibited by LLMs that hinge on subtle changes in wording and text layout between apparently similar prompts, suggesting that certain semantically similar tokens may have dramatically different topological neighborhoods in the token subspace. Linguistically, these irregularities may correspond to polysemy or homonyms (Jakubowski et al. (2020)).
@@
The presence of irregularities imply small changes in language can drive large variation in the token subspace. Moreover, if the token subspace of an LLM is, in fact, singular, hence not a manifold, irregularities can persist into the output of the LLM, perhaps unavoidably and potentially even because of its architecture. This is because the input and output of LLMs are a fixed finite set of tokens yet the internal architectures of the model assume a structure resembling a Riemannian manifold: a metrizable space that locally preserves inner products. Not accounting for singularities in the token subspace may thereby confound our understanding of the LLM’s behavior.
1.1 Contributions
Our contributions are: (1) two novel statistical tests (for manifolds and for fiber bundles; see below), (2) evidence to reject both hypotheses for several LLMs (Section 5.2), and therefore (3) the identification of otherwise unreported structure in the embeddings for these LLMs. We emphasize that while this paper demonstrates the manifold and fiber bundle tests on token embeddings, our main contribution is primarily a methodology for hypothesis testing on embedded points.
Because connected manifolds have a unique dimension, and because the dimension of fiber bundles are characterized by Theorem 1, our two tests address the following hypotheses given in Table 1.
Table 1: Hypotheses we test. We assume the token subspace has reach $\tau>0$ (see definition in Sec. 2), and estimate dimension in the ball centered at token $\psi$ with radius $r<\tau$.
Manifold test Fiber bundle test
$H_{0}$ There is a unique dimension at $\psi$ The dimension at $\psi$ in a ball of radius $r$
does not increase as $r$ increases.
$H_{1}$ There is not a unique dimension at $\psi$ The dimension at $\psi$ increases at some $r$
Both tests are implemented as Algorithm 1. The embedded reach $\tau$ defines an assumption on the curvature of the underlying space. Low reach corresponds to high curvature and vice versa (Berenfeld et al. (2022)). Classical Lagrange interpolation asserts that any finite sample of points can be fit exactly by a manifold. The resulting manifold may have extremely high curvature, which is undesirable because it results in numerical instability (Isaacson and Keller (2012)).
@@ 1.1 Contributions
To demonstrate our method, Figure 3(a)–(c) shows the application of the manifold test to several synthetic examples. In Section 5.2, we then provide empirical evidence that the token embedding function obtained from each of four different open source LLMs of moderate size—GPT2 (Radford et al. (2019)), Llemma7B (Azerbayev et al. (2024)), Mistral7B (Jiang et al. (2023)), and Pythia6.9B (Biderman et al. (2023))—cannot be a manifold due to non-constant local dimension as well as tokens (points) without an intrinsic dimension. When these tests are rejected at token $\psi$, this implies that $\psi$ is less stable under mathematical transformations than other tokens. In Section 3.2, we show the endemic nature of these instabilities by proving that context cannot resolve singularities persistently. And finally, we provide conditions (Theorem 2) for when singularities can propagate into the output of an LLM for generic transformers.
1.2 Background and related work
@@
Manifold learning takes its roots in nonlinear dimension reduction of high dimensional data into low dimensional representations of smooth manifolds. While these algorithms have played an important role in modern data science especially in domains such as cosmology (Park et al. (2024)), astrophysics (Makinen et al. (2024); Einig et al. (2023)), dynamical systems (Masuda and Kundu (2022); Parsons and Rogers (2015)), and quantum chemistry (Hemmateenejad et al. (2004)), the main utility of these methods have been for data visualization. For instance, even though UMAP provided an optimal algorithm for fitting sampled data to a manifold, the authors themselves concede that “if the global structure is of primary interest then UMAP may not be the best choice for dimensionality reduction”. A manifold learning algorithm will find some smooth structure to fit the data rather than “error out” when the data is not sampled from a manifold as it is not a hypothesis test.  Grigsby et al. (2023) showed the dimension of ReLU-Nets are inhomogenous, and so when algorithms fit a manifold to data optimally, for instance, the found manifold may have extremely high curvature or dimension. Both of these are undesirable properties in a model, and both implicitly suggest that a manifold would be a poor choice for a representation (Isaacson and Keller (2012)).
To address this concern, Fefferman et al. (2016) presented a test for whether data lie on a manifold with a specified embedded reach, but did not present the values of constants that are critical for practical implementation. More recently, Lim et al. (2023); Von Rohrscheidt and Rieck (2023) present tests for the manifold hypothesis, with promising results on low dimensional datasets. Our paper remedies both issues: we construct a practical implementation of hypothesis tests for manifolds and fiber bundles with a specified embedded reach.
@@
The token input embedding matrix defines the salient subspace of the latent space for an LLM, by specifying where the tokens are located. There is no a priori reason to suspect it is a manifold. It has already been shown that local neighborhoods of each token have salient topological structure (Rathore et al. (2023)), and one of the most basic parameters is the dimension near any given token in this space. Higher dimension at a token $\psi$ means that $\psi$ has greater density while lower dimensional tokens have less density (Jakubowski et al. (2020)). Intuitively, a densely populated neighborhood $B(\psi)$ will generically occupy a higher dimension than otherwise.
Figure 1: The histogram of local dimensions estimated near tokens in GPT2 (Robinson et al. (2024)). This histogram shows a mixture with at least three distinct peaks.
@@
Assuming that word embeddings yield manifolds, some researchers have used global dimension estimators on token input embeddings and word embeddings (Kataiwa et al. (2025); Gromov et al. (2024); Tulchinskii et al. (2023)). By using a local (not global) dimension estimator, Robinson et al. (2024) presented the first (to our knowledge) direct test of whether the token subspace is a manifold for the token input embeddings for several LLMs, although no rigorous hypothesis test was presented. The subspace of tokens has a multimodal distribution of dimension estimates (Figure 1), so unless the space is highly disconnected, it cannot be a manifold.
2 Mathematical background
A manifold is a space in which every point has a Euclidean neighborhood, and the space can be covered by taking finitely many of these neighborhoods at each point. It follows that if a manifold is connected, then the dimension of each of these Euclidean neighborhoods is the same; we call this the dimension of the manifold. Figure 3(a) shows the hollow sphere, which is an example of a $2$-dimensional manifold as a subset of $\mathbb{R}^{3}$.
Given an embedding dimension $\ell$, embedded tokens lie on a subspace $X$ of Euclidean space $\mathbb{R}^{\ell}$, with the topology and geometry inherited from $\mathbb{R}^{\ell}$. Equivalently, one can consider an embedding function $e:X\to\mathbb{R}^{\ell}$, which is one-to-one and has a full rank matrix of partial derivatives (Jacobian matrix). Embedding functions therefore preserve topological structure. The inherited information includes a metric (the Euclidean distance) and a notion of volume. The volume form on a subspace is induced by the Euclidean volume. If the subspace is a manifold that is compatible with the inherited metric and volume form, we call the subspace a Riemannian manifold. For a $2$-dimensional Riemannian manifold, the volume form is the familiar notion of area. In general, if a space is a Riemannian manifold of dimension $d$, then the volume $v$ of a ball of radius $r$ centered at any point is given by an equation of the form Gray (1974),
$v=Kr^{d}+\text{(correction terms)}r^{d+1},$ (1)
where $K>0$ is a constant depending on $d$. If the correction terms are large, the manifold is said to have high curvature.
The definition of a manifold is global: if even one point fails to have a Euclidean neighborhood, then the space cannot be a manifold, and such a point is called a singularity. One commonly encountered singularity is a boundary, in which a point has a neighborhood that is topologically equivalent to a half space. Spaces whose only singularities are boundaries are called manifolds with boundary. For instance, the solid ball in $\mathbb{R}^{3}$ is a manifold with boundary. Its set of singularities is the hollow $2$-dimensional sphere.
Classifying the types of singularities up to topological equivalence of their neighborhoods is in general an open problem, though some types are common. Figure 3(b) shows a space with cusp and boundary singularities. Figure 3(c) shows a singularity associated with a change in dimension.
In order to handle singularities, we also consider a generalization of a manifold called a fibered manifold with boundary (fiber bundle for short). This is a space in which every point has a neighborhood that can be written as a Cartesian product $B\times F$, where $B$ is a manifold and $F$ is a manifold with boundary. We call $B$ the base space and $F$ the fiber space.
Curvature (the correction terms in Equation (1)) does not generalize cleanly to spaces with singularities. Instead, the embedded reach plays a similar role in what follows. The embedded reach (simply reach) for a space $X\subseteq\mathbb{R}^{\ell}$ is the largest distance $\tau\geq 0$ that a point $y\in\mathbb{R}^{\ell}$ can be from $X$, yet still have a unique closest point on $X$. For convex subspaces, $\tau=\infty$. Conversely, spaces with a cusp (like Figure 3(b)), have $\tau=0$. Manifolds with small reach have high curvature, and vice versa (Berenfeld et al. (2022)).
3 Theoretical insights
@@
Theorem 1 asserts that for a fiber bundle, the number of tokens within a ball of a given radius $r$ centered at a fixed token $\psi$ is a piecewise log-log-linear function of radius, in which the slopes correspond to dimension, and the slopes cannot increase as $r$ increases.
3.1 Volume versus radius for fiber bundles
@@
The learned token embedding vectors in an LLM each have the same number of elements, which we refer to as $\ell$. Thus, the explicit representation of the token subspace $T$ used by an LLM arises by an embedding function $e:T\to\mathbb{R}^{\ell}$ into the latent space $\mathbb{R}^{\ell}$. Since embeddings are one-to-one, each token $\psi\in T$ corresponds to a unique element $e(\psi)\in\mathbb{R}^{\ell}$ in the latent space. With a slight abuse of notation, we can write $\psi$ for both.
@@
Suppose that $T$ is a compact, finite-dimensional Riemannian manifold with boundary, with a volume form $v$ satisfying $v(T)<\infty$, and let $p:T\to S$ be a fiber bundle. If $e:T\to\mathbb{R}^{\ell}$ is a smooth embedding with reach $\tau$, then there is a function $\rho:e(T)\to[0,\tau]$ such that if $\psi\in e(T)$, the induced volume $(e_{*}v)$ in $\mathbb{R}^{\ell}$ satisfies
[…]
@@
See Section A.5 for full details; a sketch follows. Taking the logarithm of both sides of Equation (1) yields a linear equation, in which $d$ is the slope.
For the case of the fiber bundle, Equation (1) must be modified to handle the case where the ball of radius $r$ “sticks out” of the space. This causes a reduction in volume, hence a smaller slope. ∎
The probability distribution of tokens in $T$ can be modeled by the volume form $v$. If a fiber bundle is the correct representation of the space of tokens, Theorem 1 characterizes the resulting probability distribution in $\mathbb{R}^{\ell}$ using parameters (the exponents in Equation (2)) that can be estimated from the token input embedding. These parameters are bounded by the dimensions of $S$ and $T$. In Algorithm 1, Monte Carlo estimates of volume are used, simply by counting tokens within a given radius. Although the convergence of Monte Carlo estimates is a strong assumption, it is also extremely well-established in the literature.
Figure 2: Interpretation of the manifold and fiber bundle tests: the manifold test detects any slope changes in the log volume versus log radius curves, while the fiber bundle test only detects slope increases.
In short, if we take the logarithm of both the volume and radius in Equation (2), we obtain a linear curve for a manifold. Abrupt slope changes imply that the space is not a manifold. For a fiber bundle, we obtain a piecewise linear curve, in which the slopes are given by $\dim S<\dim T$, and the slopes must decrease through a discontinuity as the radius increases, as shown in Figure 2.
3.2 Context does not persistently resolve singularities
The remaining question is whether singularities can persist into the output of the LLM. Can the user elicit the instabilities caused by singularities? We provide sufficient conditions for unstable outputs:
Theorem 2.
Let $Z$ be a $d$-dimensional bounding manifold [^1] Since $T\subseteq\mathbb{R}^{\ell}$, it is always contained within such a bounding manifold $Z\subseteq\mathbb{R}^{\ell}$. for the token subspace, such that $T\subseteq Z$. Consider an LLM with a context window of size $w$, in which the latent space of tokens is $\mathbb{R}^{\ell}$, and we collect $m$ tokens as output from this LLM.
Suppose the following, (enough tokens are collected from the response) $m>\frac{2wd}{\ell}$, but (the context window is longer than the number of tokens we collected) $w\geq m$. Under these conditions, a generic set of transformers yields a topological embedding of $T^{w}=T\times\dotsb T$ into the output of the LLM.